# Learning to Adapt SAM for Segmenting Cross-domain Point Clouds

### v3

@@ 4.3 Experimental Results and Comparison
Tab. 1 and Tab. 2 show the experimental results and performance comparison with previous UDA methods for 3D segmentation under the setup introduced in Sec. 4.1. Each experiment contains two reference methods in common, a baseline model named Source only trained only on the source domain and an upper-bound model named Oracle trained only on the target data with annotations. Tab. 1 focuses on four domain shift scenarios introduced by  [20] and comparison with these multi-modal methods based on xMUDA such as  [32] and  [5]. Among them, MM2D3D fully exploits the complementarity of image and point cloud and proposes to feed two modalities to both branches, achieving better performance. Our method outperforms it by $+6.8\%$ (USA $\rightarrow$ Singapore), $+0.3\%$ (Day $\rightarrow$ Night), $+14.6\%$ (v.KITTI $\rightarrow$ Sem.KITTI), $+6.0\%$ (A2D2 $\rightarrow$ Sem.KITTI) respectively, because our method aligns all the features into a unified feature space with the guidance of SAM instead of simply aligning features from image and point cloud in 2D and 3D network. Tab. 2 focuses on the scenarios where not all the point clouds are visible in the images and we re-implement three methods by their official codes.  [30] uses the prediction from the pre-trained model as pseudo labels for unlabelled data to retrain this model, which is widely used in UDA methods.  [37] trains a 3D network with only the utilization of a point cloud, which generates new intermediate domains through a mixing scene-level strategy to mitigate domain discrepancies. MM2D3d is the SOTA multi-modal method as described above, but it needs all the points visible in the image for the best performance. Our method surpasses them by at least $+17.9\%$ (nuScenes $\rightarrow$ Sem.KITTI), $+11.0\%$ (Sem.KITTI $\rightarrow$ nuScenes), $+13.0\%$ (nuScenes $\rightarrow$ Waymo), $+14.7\%$ (Waymo $\rightarrow$ nuScenes) respectively by a large margin, since hybrid feature augmentation can provide more intermediate domains and SAM-guided feature alignment can help map the whole point cloud into the unified feature space.
@@
According to the results above, our method outperforms others by a large margin, attributed to the full utilization of the general knowledge provided by SAM. To further prove that the achievement is due to not only SAM but also our novel paradigm, we also use SAM for current SOTA multi-modal UDA work and get results in the row of ”MM2D3D*”. Because it trains both 2D and 3D networks simultaneously, under its original framework, we can only refine the supervision signals of 2D network using the instance mask output of SAM. As seen from the results, SAM can improve its performance but very limited. Our method not only uses instance masks but also makes full use of the general features extracted by SAM, ensuring the superiority of our method. Qualitative results are shown in Fig. 4, where predictions in the ellipses demonstrate our method can provide correct and more fine-grained segmentation.
Figure 4: Visualization results of the domain adaptation from nuScenes to SemanticKITTI.
@@
Setting Baseline SAM-guided Feature Align. Hybrid Feature Aug. Pseudo Label mIoU
 Scene Instance
@@ 4.4 Ablation Study
Effectiveness of Hybrid Feature Augmentation For the detailed ablation of feature alignment, we adopt hybrid strategies for diverse data with distinct feature patterns, which not only mix up points at the scene level in polar-based, range-based, and laser-based ways but also at the instance level with the help of instance mask output by SAM. Random selection in all these point mix-up ways forms this feature augmentation. Setting (4), (5), (6) in Tab. 3 shows that both mix-up methods can help feature alignment with more distinct features but the hybrid strategy raises the best performance. More ablations are in the Appendix.
Effectiveness of Vision Foundation Model Our insight design is using VFM to provide a unified feature space to align point features. To verify the idea is reliable, we randomly take amounts of data from the source and target datasets and calculate the difference of feature distributions from source and target domain by KL divergence. As Fig. 5 shows, the feature distribution differences after mapping to SAM feature space are much smaller than by using original 3D point features, making our feature alignment much easier. Moreover, we also extend our method to other vision foundation models, such as InternImage [40], serving for image-based tasks. Specifically, we replace the SAM-based image encoder with InternImage to guide the feature alignment in a similar manner. Compared with the baseline (mIoU=27.7), the performance of using InternImage is mIoU=36.9. A consistent performance gain can be obtained, which also verifies and validates our insight, i.e., the generic feature space of the VFM can ease the feature alignment.
Figure 5: Visualization of the feature distance across different datasets, where smaller values indicate more similar distribution.
@@ 4.5 More Challenging Tasks
Since we achieve the purpose of mapping data from different domains into a unified feature space, the extracted feature can be used for some more challenging tasks. We show some extension results of our method in Tab. 4. The top subtable shows the results of UDA for panoptic segmentation, a more challenging task requiring instance-level predictions. With more accurate and fine-grained semantic prediction, our method achieves promising results. The down subtable shows the results of domain generalization, in which target data only can be used for testing. In this subtable, models are trained with nuScenes and SemanticKITTI, then evaluated with A2D2 dataset. With the ability of stronger data-to-feature mapping, our method outperforms the current SOTA method [27]. In the future, we seek to explore the potential of our method on more tasks, such as 3D detection.
Table 4: Extension on more challenging tasks, such as UDA for Panoptic Segmentation(up) and Domain Generalization(down).
Task Method PQ ${}^{\dagger}$ RQ SQ mIoU
@@
Method baseline xMUDA [20] Dual-Cross [26] BEV-DG [27] Ours
mIoU 45.0 44.9 41.3 55.1 57.2
@@ 5 Conclusion
6 Acknowledgements
This work was supported by NSFC (No.62206173), Natural Science Foundation of Shanghai (No.22dz1201900), MoE Key Laboratory of Intelligent Perception and Human-Machine Collaboration (ShanghaiTech University), Shanghai Frontiers Science Center of Human-centered Artificial Intelligence (ShangHAI), Shanghai Engineering Research Center of Intelligent Vision and Imaging.
Supplementary Material
In this section, We provide more supplementary material to support the findings and observations drawn in the main body of this paper.
A. Dataset Details
We conduct extensive experiments under several domain shift scenarios with five well-known large-scale datasets, including nuScenes, SemanticKITTI, VirtualKITTI, A2D2, and Waymo. All of these datasets provide point clouds and corresponding images captured by distinct devices resulting in different data representations, thus we will give more detailed information about them. (1) nuScenes contains 1000 driving scenes with 20 seconds for each scene, taken at 2Hz. The scenes are split into train (28,130 keyframes), validation (6,019 keyframes), and corresponding point-wise 3D semantic labels provided by nuScenes-Lidarseg. (2) SemanticKITTI features a large-angle front camera and a 64-layer LiDAR and the captured data from scenes 0, 1, 2, 3, 4, 5, 6, 7, 9, and 10 are used for training while scenes 8 as validation at most experiments. (3) VirtualKITTI consists of 5 driving scenes created with the Unity game engine. VirtualKITTI does not simulate LiDAR but rather provides a dense depth map with semantic labels, so we use the 2D-to-3D projecting to generate a point cloud from the depth map. (4) A2D2 consists of 20 drives corresponding to 28,637 frames. The point cloud comes from three 16-layer front LiDARs (left, center, right), and the semantic labeling was carried out in the 2D image. (5) Waymo offers 2,860 temporal sequences captured by five cameras and one LiDAR in three different geographical locations, leading to a total of 100k labeled data, making it larger than existing datasets that offer point-wise segmentation labels. We visualize these five datasets with one sample to show the data domain difference in Fig. 6.
Table 5: Effect of different point mix-up strategies for the hybrid feature augmentation.
Scene-level Feature Augmentation Instance-level Feature Augmentation mIoU
Polar-based Range-based Laser-based
$\checkmark$ 37.0
$\checkmark$ 38.5
$\checkmark$ 37.6
$\checkmark$ 40.1
$\checkmark$ $\checkmark$ $\checkmark$ 39.0
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 44.0
Figure 6: Visualization of each used dataset. From left to right, the figure shows the point cloud, one image corresponding to the point cloud, and the projection of the point cloud on the image.
B. More Experimental Data
We conduct further ablation on the same setting in the Ablation Study for hybrid feature augmentation to thoroughly analyze its effect. We take all the point mix-up strategies used in it apart to conduct separate experiments as shown in Tab. 5 and these results show that all of them contribute to our method. So we combine them with random selection to form the hybrid feature augmentation. This comprehensive approach boosts the performance since it can provide different data with distinct feature patterns as much as possible for SAM-guided feature alignment. Additionally, we found that this approach can be used in other multi-modal methods as a normal data augmentation method. Taking MM2D3D as an example, we add this module into it and Tab. 6 shows that our feature augmentation approach also improves its performance. However, this feature augmentation approach contributes to our method in a better way because of our effective SAM-guided feature alignment.
Table 6: Generalization of feature augmentation for other multi-modal UDA methods.
Source only MM2D3D MM2D3D w/ feature augmentation Ours w/o feature augmentation Ours
27.7 30.4 33.2 34.0 48.5
C. More Qualitative Results
Except for the visualization results of the domain adaptation from nuScenes to SemanticKITII shown in the main paper, we provide additional qualitative results in Fig. 7 representing other introduced domain shift scenarios in the experiment including SemanticKITII-to-nuScenes, nuScenes-to-Waymo, and Waymo-to-nuScenes. Our method has more accurate predictions for the cars, sidewalk, terrain, etc.
Figure 7: Visualization results of the domain adaptation from more domain shift scenarios, including (1)SemanticKITII-to-nuScenes, (2) nuScenes-to-Waymo, and (3) Waymo-to-nuScenes.

### v4

@@ 4.3 Experimental Results and Comparison
Tab. 1 and Tab. 2 show the experimental results and performance comparison with previous UDA methods for 3D segmentation under the setup introduced in Sec. 4.1. Each experiment contains two reference methods in common, a baseline model named Source only trained only on the source domain and an upper-bound model named Oracle trained only on the target data with annotations. Tab. 1 focuses on four domain shift scenarios introduced by xMUDA [20] and comparison with these multi-modal methods based on xMUDA such as DsCML [32] and MM2D3D [5]. Among them, MM2D3D fully exploits the complementarity of image and point cloud and proposes to feed two modalities to both branches, achieving better performance. Our method outperforms it by $+6.8\%$ (USA $\rightarrow$ Singapore), $+0.3\%$ (Day $\rightarrow$ Night), $+14.6\%$ (v.KITTI $\rightarrow$ Sem.KITTI), $+6.0\%$ (A2D2 $\rightarrow$ Sem.KITTI) respectively, because our method aligns all the features into a unified feature space with the guidance of SAM instead of simply aligning features from image and point cloud in 2D and 3D network. Tab. 2 focuses on the scenarios where not all the point clouds are visible in the images and we re-implement three methods by their official codes. PL [30] uses the prediction from the pre-trained model as pseudo labels for unlabelled data to retrain this model, which is widely used in UDA methods. CosMix [37] trains a 3D network with only the utilization of a point cloud, which generates new intermediate domains through a mixing scene-level strategy to mitigate domain discrepancies. MM2D3d is the SOTA multi-modal method as described above, but it needs all the points visible in the image for the best performance. Our method surpasses them by at least $+17.9\%$ (nuScenes $\rightarrow$ Sem.KITTI), $+11.0\%$ (Sem.KITTI $\rightarrow$ nuScenes), $+13.0\%$ (nuScenes $\rightarrow$ Waymo), $+14.7\%$ (Waymo $\rightarrow$ nuScenes) respectively by a large margin, since hybrid feature augmentation can provide more intermediate domains and SAM-guided feature alignment can help map the whole point cloud into the unified feature space.
@@
According to the results above, our method outperforms others by a large margin, attributed to the full utilization of the general knowledge provided by SAM. To further prove that the achievement is due to not only SAM but also our novel paradigm, we also use SAM for current SOTA multi-modal UDA work and get results in the row of "MM2D3D*". Because it trains both 2D and 3D networks simultaneously, under its original framework, we can only refine the supervision signals of 2D network using the instance mask output of SAM. As seen from the results, SAM can improve its performance but very limited. Our method not only uses instance masks but also makes full use of the general features extracted by SAM, ensuring the superiority of our method. Qualitative results are shown in Fig. 4, where predictions in the ellipses demonstrate that source-only and MM2D3D models often infer wrong and mingling results, especially for the person category, while our method can provide correct and more fine-grained segmentation. More qualitative results are in the Appendix.
Figure 4: Visualization of the domain adaptation from nuScenes to SemanticKITTI.
@@
Setting Baseline SAM-guided Feature Alignment Hybrid Feature Augmentation Pseudo Label mIoU
 Scene-level Instance-level
@@ 4.4 Ablation Study
Effectiveness of Hybrid Feature Augmentation For the detailed ablation of feature alignment, we adopt hybrid strategies for diverse data with distinct feature patterns, which not only mix up points at the scene level in polar-based, range-based, and laser-based ways but also at the instance level with the help of instance mask output by SAM. Random selection in all these point mix-up ways forms this feature augmentation. Setting (4), (5), (6) in Tab. 3 shows that both mix-up methods can help feature alignment with more distinct features but the hybrid strategy raises the best performance. More ablations are in the Appendix. Moreover, since masks generated by SAM do not contain semantic labels, we conduct an additional experiment to prove the validity of instance-level augmentation by replacing the generated SAM instance masks with the ground truth semantic masks under the setting (6). Compared with the original result (mIoU=44.0), the performance of using ground truth semantic masks is mIoU=42.7, demonstrating that although masks generated by SAM cannot be identical to the ground truth, the contained semantic information is consistent, and the randomness of our augmentation further improve the performance.
Table 4: The effect of image-point coverage. Different numbers represent the number of used pictures of nuScenes.
Covered Images for Alignment baseline 2 4 6
nuScenes $\rightarrow$ Sem.KITTI 27.7 43.9 45.6 48.5
Sem.KITTI $\rightarrow$ nuScenes 28.1 38.4 40.6 42.9
Effect of Different Point-to-Pixel Coverage for Alignment. For point clouds not covered by images, we do not calculate feature alignment loss and solely calculate segmentation loss with ground truth or pseudo label. When conducting the experiments, we used all available images to ensure that as many point cloud data as possible could find corresponding features on the pictures for SAM-guided feature alignment. We also conduct additional experiments to demonstrate the impact of the number of available images for feature alignment, as shown in Tab.4. The results indicate that a greater number of available images correlates with improved experimental outcomes. Importantly, as long as SAM-guided feature alignment is achievable, the performance does not significantly degrade even with limited coverage.
Effect of Vision Foundation Model We also extend our method to other vision foundation models, such as InternImage [41], serving for image-based tasks. Specifically, we replace the SAM-based image encoder with InternImage to guide the feature alignment in a similar manner. Compared with the baseline (mIoU=27.7), the performance of using InternImage is mIoU=36.9. A consistent performance gain can be obtained, which also verifies and validates our insight, i.e., the generic feature space of the VFM can ease the feature alignment.
@@ 4.5 More Challenging Tasks
Since we achieve the purpose of mapping data from different domains into a unified feature space, the extracted feature can be used for some more challenging tasks. We show some extension results of our method in Tab. 5. The left subtable shows the results of UDA for panoptic segmentation, a more challenging task requiring instance-level predictions. With more accurate and fine-grained semantic prediction, our method achieves promising results. The right subtable shows the results of domain generalization, in which target data only can be used for testing. In this subtable, models are trained with nuScenes and SemanticKITTI and then evaluated with A2D2 dataset. With the ability of stronger data-to-feature mapping, our method outperforms the current SOTA method [26]. In the future, we seek to explore the potential of our method on more tasks, such as 3D detection.
Table 5: Extension on more challenging tasks, such as UDA for Panoptic Segmentation(left) and Domain Generalization(right), where N, S, A represent nuScenes, SemanticKITTI and A2D2 dataset.
Task Method PQ PQ RQ SQ mIoU
@@
Method N,S $\rightarrow$ A
Baseline $45.0$
xMUDA [20] $44.9$
Dual-Cross [27] $41.3$
BEV-DG [26] $55.1$
Ours $\mathbf{57.2}$